# Pauli blocking: probing beyond-mean-field effects in neutrino flavor evolution

### v1

@@ Abstract
Neutrino quantum kinetics in dense astrophysical environments is investigated relying on the mean-field approximation. However, it remains to be understood whether mean-field corrections could hinder flavor instabilities that are otherwise foreseen. In this paper, we heuristically explore whether beyond-mean-field effects due to neutrino degeneracy can affect the flavor conversion physics. We find that these corrections shift the stability regions for a suite of (anti)neutrino distributions; a configuration of angular distributions that is stable in the mean-field case can become unstable, or the flavor conversion of previously unstable ensembles can be damped. Our work should serve as a motivation for further understanding the limitations of the mean-field treatment.
@@ I Introduction
Fast flavor conversion is expected to occur in the neutrino decoupling regions of core-collapse supernovae and neutron-star merger remnants, with potential implications for the neutrino-driven explosion mechanism and the nucleosynthesis of elements heavier than iron [25, 26, 27, 28, 29, 30, 31, 32]. However, our current understanding of collective flavor conversion is based on the solution of neutrino quantum kinetic equations that employ the mean-field approximation [5, 33]. The validity of the mean-field approach and the nature of the resulting phenomenology are the subject of investigation. For example, ongoing work attempts to look for many-body corrections to the mean-field treatment (cf., e.g., Ref. [34] and references therein), but tractable systems are too idealized to give an insight on any eventual implications on astrophysical sources [35, 36].
@@
Many-body treatments allow for correlations between particles with different momenta, which are explicitly forbidden at the mean-field level. Such correlations could arise due to entanglement or close to the core of compact astrophysical sources, where the baryon density reaches nuclear levels, and (anti)neutrinos are degenerate and undergo frequent collisions with matter: see the sketch in Fig. 1. Hence, there is reason to ask whether they could lead to novel neutrino flavor instabilities. The region where beyond-mean-field corrections due to neutrino degeneracy could affect flavor evolution can be constrained a priori. At nuclear densities, neutrinos of all flavors are in thermal equilibrium with matter because of frequent interactions with the background. The chemical potential for the heavy-lepton flavors ($\nu_{x}$ and $\bar{\nu}_{x}$) is negligible, and for electron (anti)neutrinos, $\nu_{e}$ ($\bar{\nu}_{e}$), it is positive (negative). This leads to the following hierarchy between the number densities of the different neutrino species: $n_{\nu_{e}}>n_{\nu_{x}}>n_{\bar{\nu}_{e}}$. Hence, as long as neutrinos are trapped, flavor instabilities are disfavored. At larger distances from the source core, as the baryon density decreases and neutrino decoupling occurs, $n_{\nu_{x}}<n_{\nu_{e}},n_{\bar{\nu}_{e}}$; in this region, the (anti)neutrino angular distributions become forward peaked and ELN crossings develop [37, 38, 39]. This means that fast flavor instabilities can occur, but momentum correlations are unlikely to have an effect due to low neutrino density. If beyond-mean-field corrections caused by neutrino degeneracy should have an impact on flavor evolution, their effect should originate in the region of transition between the trapping and free-streaming regimes.
@@
In this paper, we explore the impact of beyond-mean-field corrections due to neutrino degeneracy on neutrino quantum kinetics and, in particular, whether flavor instabilities can arise from or be suppressed by these effects. This manuscript is structured as follows. In Sec. II, we provide an overview of the neutrino equations of motion in the presence of neutrino self-interactions and introduce a heuristic prescription aiming to gauge beyond-mean-field effects linked to neutrino degeneracy. In Sec. III, we linearize the neutrino equations of motion and, relying on a toy model, compute the growth rate of flavor instabilities for a set of ELN angular distributions, within the single-energy approximation. In Sec. IV, we present the solutions of the neutrino equations of motion for three benchmark ELN distributions. To understand the flavor evolution in the non-linear regime in the presence of neutrino degeneracy, we compare our results with the well-known case of the fast flavor pendulum [16, 21]. In Sec. V, we relax the single-energy approximation and assess whether the energy dependence further affects the development of flavor instabilities. Finally, we discuss our findings and conclude in Sec. VI.
@@ II Neutrino equations of motion
$\rho_{\mathbf{r},\mathbf{p}}$ $=\begin{pmatrix}\rho_{ee;\mathbf{r},\mathbf{p}}&\rho_{ex;\mathbf{r},\mathbf{p}}\\ \rho_{ex;\mathbf{r},\mathbf{p}}^{*}&\rho_{xx;\mathbf{r},\mathbf{p}}\end{pmatrix}\ ,$ (1a)
$\bar{\rho}_{\mathbf{r},\mathbf{p}}$ $=\begin{pmatrix}\bar{\rho}_{ee;\mathbf{r},\mathbf{p}}&\bar{\rho}_{ex;\mathbf{r},\mathbf{p}}\\ \bar{\rho}_{ex;\mathbf{r},\mathbf{p}}^{*}&\bar{\rho}_{xx;\mathbf{r},\mathbf{p}}\end{pmatrix}\ ,$ (1b)
where $\rho_{ll^{\prime};\mathbf{r},\mathbf{p}}$ and $\bar{\rho}_{ll^{\prime};\mathbf{r},\mathbf{p}}$ denote the number density ($l=l^{\prime}$) and flavor coherence ($l\neq l^{\prime}$). These quantities represent an average over a certain volume (momentum spread) around the position $\mathbf{r}$ (momentum $\mathbf{p}$). Equation (1) carries the assumption that inhomogeneities only occur at large scales, e.g. due to temperature gradients in the medium [5, 40, 33]. Short-range fluctuations would be associated with large momentum uncertainties, which would make it impossible to specify $\mathbf{p}$. Moreover, in a strongly inhomogeneous system, different momenta would be correlated, so $\rho_{\mathbf{r},\mathbf{p}}$ and $\bar{\rho}_{\mathbf{r},\mathbf{p}}$ cannot fully describe the ensemble [41]. In this section, we first introduce the equations of motion of neutrinos within the mean-field approximation, and then modify them to take into account effects linked to neutrino degeneracy not included in the mean-field approach.
In the following, we assume that the system is fully axisymmetric and normalize the density matrices such that $(2\pi)^{-1}\int\mathrm{d}^{3}p\mathrm{Tr}(\rho_{\mathbf{p}})=1$ and $(2\pi)^{-1}\int\mathrm{d}^{3}p\mathrm{Tr}(\bar{\rho}_{\mathbf{p}})=\bar{n}_{\nu}/n_{\nu}$, where $n_{\nu}$ ($\bar{n}_{\nu}$) is the total (anti)neutrino number density. A factor $2\pi$ from the integral over the azimuthal angle is absorbed into the definition of $\rho_{\mathbf{r,p}}$ and $\bar{\rho}_{\mathbf{r,p}}$, and the momentum distribution can be parametrized in terms of energy, $E=|\mathbf{p}|c$, and radial velocity $vc$.
@@ II.1 Equations of motion within the mean-field approximation
The mean-field approximation assumes that, locally, all (anti)neutrinos can be described as plane waves. This approach prevents any correlation between neutrinos with different momenta. Neutrinos with the same flavor are indistinguishable, so their quantum state has to be antisymmetric to particle exchange, and momentum correlations should appear naturally. Hence, $\rho_{\mathbf{r},\mathbf{p}}$ and $\bar{\rho}_{\mathbf{r},\mathbf{p}}$ do not fully describe a quantum state; they determine a phase-space distribution that averages over different particles.
Within this framework, the evolution of the density matrices is driven by coherent forward scattering, i.e. refraction. In particular, if a neutrino can interact with many identical neutrinos, the transition amplitudes of all allowed processes add up coherently. In a homogeneous medium, the interference terms enhance elastic processes such as:
${{\nu}}_{l}(\mathbf{p})+{\nu}_{l^{\prime}}(\mathbf{q})$ $\to{\nu}_{l}(\mathbf{q})+{\nu}_{l^{\prime}}(\mathbf{p})\,,$ (2a)
$\nu_{l}(\mathbf{p})+\bar{\nu}_{l}(\mathbf{q})$ $\to\nu_{l^{\prime}}(\mathbf{p})+\bar{\nu}_{l^{\prime}}(\mathbf{q})\,.$ (2b)
These interactions provide the mechanism behind collective flavor conversion, as they can amplify the flavor coherence generated by the neutrino mass difference. Their effect on a purely homogeneous (anti)neutrino ensemble ($\rho_{\mathbf{r,p}}\to\rho_{\mathbf{p}}$ and $\bar{\rho}_{\mathbf{r,p}}\to\bar{\rho}_{\mathbf{p}}$) is described by the following Hamiltonian:
@@
where ${D}_{n}\equiv\int\mathrm{d}^{3}p^{\prime}(2\pi E^{\prime 2})^{-1}L_{n}(v^{\prime})({\rho}_{\mathbf{p^{\prime}}}-\bar{\rho}_{\mathbf{p^{\prime}}})$. Here, $L_{n}$ are the Legendre polynomials, normalized such that $\int_{-1}^{+1}\mathrm{d}v^{\prime}L_{n}(v^{\prime})L_{m}(v^{\prime})=(n+\frac{1}{2})^{-1}$, and $v^{\prime}c$ is the radial (anti)neutrino velocity.
The density matrices do not describe an antisymmetric quantum state, but Eqs. (4) cannot violate the exclusion principle. Even if the mean-field framework is extended to include incoherent processes (collisions), which can change the number of particles, the scattering rates tend to 00 as the (anti)neutrino occupation numbers increase [5]. Hence, no momentum mode can have an occupation number greater than $1$. However, it is important to question how neglecting momentum correlations shapes our understanding of flavor conversion. For example, it is usually assumed that the system starts as a mixture of pure flavor eigenstates (i.e. the density matrices are diagonal). This is motivated when describing astrophysical environments, where (anti)neutrinos are emitted as highly localized wavepackets with a specific flavor: if these functions do not overlap, then the system is fully incoherent. On the other hand, two plane waves with the same momentum always interfere.
@@ II.2 Beyond-mean-field effects linked to neutrino degeneracy
If the ensemble is inhomogeneous, the momenta in Eq. (2) are approximate quantities. The interacting particles do not exactly swap their momenta, nor are they scattered precisely in the forward direction. The initial and final momenta are indistinguishable within their uncertainties. In particular, this means that every neutrino contributing to the scattering amplitude can recoil to some extent. This momentum exchange cannot be captured by the mean-field approximation and could modify the development of coherence (see Fig. 2): neutrinos cannot recoil into states that are already occupied, so certain coherent processes could be Pauli blocked, if the neutrino ensemble is degenerate.
Figure 2: Neutrino-neutrino interactions in the mean-field approximation (left) vs. beyond-mean-field case (right). In both cases, the incoming $\nu_{e}$ (in blue) is in a flavor eigenstate, and a small seed of flavor coherence is generated by vacuum oscillations (the incoming $\nu_{1}$ and the outgoing neutrinos are represented as superpositions of $\nu_{e}$ and $\nu_{x}$ eigenstates, with $\nu_{x}$ being plotted in orange). In the mean-field scenario, no momentum is exchanged; the neutrino scatters in the forward direction and its flavor coherence grows as a result. In the beyond-mean-field picture, a small momentum exchange takes place, so the incoming $\nu_{e}$ is scattered onto a region where there already are $\nu_{x}$’s. The latter do not participate in the interaction, but alter the development of flavor coherence.
@@
Working out the form of this non-Hermitian Hamiltonian exceeds the scope of this paper. We attempt to assess whether neutrino degeneracy could lead to beyond-mean-field corrections that would affect flavor instabilities. The occurrence of fast instabilities depends on the presence of a crossing in the ELN angular distribution [15, 43]. However, this requisite can be relaxed, e.g., by effects such as vacuum mixing. The mass difference term in the neutrino dispersion relation can be absorbed into the definition of an effective ELN angular distribution, with stability being determined by crossings in the latter [44]. Likewise, we explore the possibility that momentum correlations in the self-interaction Hamiltonian can be modeled with effective density matrices, $\tilde{\rho}_{\mathbf{p}}$ and $\tilde{\bar{\rho}}_{\mathbf{p}}$.
In order to motivate the form of these effective density matrices, we note that, if all quantum states are filled, no net flavor conversion should take place. Thus, it is reasonable to assume that, within a single-particle framework, any corrections to Eqs. (4) should scale with the (anti)neutrino occupation numbers. Since the expected effect is a flavor-dependent suppression of coherent scattering, we propose the following ansatz:
@@
where $\mu_{\nu_{l}}$ is the chemical potential of neutrinos with flavor $l$, and $T$ is the local temperature. The $1/2$ exponent in Eqs. (6) is motivated by the fact that, while neutrino cross sections are suppressed by $\prod_{i}(1-B_{i})$, it is the scattering amplitude that enters the calculation of the neutrino refractive index.
In the rest of the paper, we investigate the effect of such beyond-mean-field corrections due to neutrino degeneracy. We stress that while the mean-field equations presented in Ref. [5] do take into account Pauli blocking effects, momentum correlations are not considered in the mean-field approach.

### v2

@@ Abstract
Neutrino quantum kinetics in dense astrophysical environments is investigated relying on the mean-field approximation. In this paper, we heuristically explore whether beyond-mean-field effects due to neutrino degeneracy could hinder flavor instabilities that are otherwise foreseen. Our results show that these corrections shift the stability regions for a suite of (anti)neutrino ensembles: the flavor conversion of previously unstable distributions can be damped, but angular distributions that are stable in the mean-field case can also become unstable. Our work should serve as a motivation to further investigate the limitations of the mean-field treatment.
@@ I Introduction
Fast flavor conversion is expected to occur in the neutrino decoupling regions of core-collapse supernovae and neutron-star merger remnants, with potential implications for the neutrino-driven explosion mechanism and the nucleosynthesis of elements heavier than iron [25, 26, 27, 28, 29, 30, 31, 32]. However, our current understanding of collective flavor conversion is based on solutions of neutrino quantum kinetic equations that employ the mean-field approximation [5, 33]. The validity of the mean-field approach and the nature of the resulting phenomenology are the subject of ongoing investigation. For example, recent work attempts to look for many-body corrections to the mean-field treatment (cf., e.g., Ref. [34] and references therein), but tractable systems are too idealized to give an insight on any eventual implications on astrophysical sources [35, 36].
@@
The mean-field neutrino-neutrino Hamiltonian describes the interaction between a neutrino field, which represents neutrinos with a certain momentum, and an external potential generated by other neutrinos. In this picture, many-body correlations between particles with different momenta (“momentum correlations”) are explicitly forbidden. As a result, all neutrinos are in momentum eigenstates and scattering is only allowed in the form of a momentum swap. Momentum correlations would be most significant close to the core of compact astrophysical sources, where neutrinos are very degenerate (see the sketch in Fig. 1). We argue that said correlations could, in fact, depend on the shape of the neutrino spectrum, and that degeneracy could therefore lead to novel effects, e.g. flavor instabilities.
@@
The region where beyond-mean-field corrections due to neutrino degeneracy could affect flavor evolution can be constrained a priori. At nuclear densities, neutrinos of all flavors are in thermal equilibrium with matter because of frequent interactions with the background. The chemical potential for the heavy-lepton flavors ($\nu_{x}$ and $\bar{\nu}_{x}$) is negligible, and for electron (anti)neutrinos, $\nu_{e}$ ($\bar{\nu}_{e}$), it is positive (negative). This leads to the following hierarchy between the number densities of the different neutrino species: $n_{\nu_{e}}>n_{\nu_{x}}>n_{\bar{\nu}_{e}}$. Hence, as long as neutrinos are trapped, flavor instabilities are disfavored. At larger distances from the source core, as the baryon density decreases and neutrino decoupling occurs, $n_{\nu_{x}}<n_{\nu_{e}},n_{\bar{\nu}_{e}}$; in this region, the (anti)neutrino angular distributions become forward peaked and ELN crossings develop [37, 38, 39]. This means that fast flavor instabilities can occur, but momentum correlations are unlikely to have an effect due to low neutrino density. If beyond-mean-field corrections caused by neutrino degeneracy should have an impact on flavor evolution, their effect should originate in the region of transition between the trapping and free-streaming regimes.
In this paper, we explore the impact of beyond-mean-field corrections due to neutrino degeneracy on neutrino quantum kinetics and, in particular, whether flavor instabilities can arise from or be suppressed by these effects. We stress that we do not aim to provide a self-consistent formalism to describe this many-body phenomenon. Our goal is to test whether many-body effects caused by neutrino degeneracy could modify the development of flavor instabilities. Since the many-body equations of motion of neutrinos embedding this process are yet to be developed, we heuristically modify the single-particle equations of motion to assess whether dedicated formal work on the development of new neutrino equations of motion is justified. This work should serve as a testbed to eventually motivate further work to extend the commonly adopted mean-field approach. The corrections to the self-interaction Hamiltonian that we introduce are inspired by the influence of the neutrino chemical potential on incoherent scattering. The neutrino emission rates are suppressed at energies $E\ll\mu_{\nu_{l}}$, because (some of) the final states in the neutrino phase space are already occupied: this effect is known as Pauli blocking.
@@
The manuscript is structured as follows. In Sec. II, we provide an overview of the neutrino equations of motion in the presence of neutrino self-interactions and introduce a prescription aiming to gauge beyond-mean-field effects linked to neutrino degeneracy. In Sec. III, we linearize the neutrino equations of motion and, relying on a toy model, compute the growth rate of flavor instabilities for a set of ELN angular distributions, within the single-energy approximation. In Sec. IV, we present the solutions of the neutrino equations of motion for three benchmark ELN distributions. To understand the flavor evolution in the non-linear regime in the presence of neutrino degeneracy, we compare our results with the well-known case of the fast flavor pendulum [16, 21]. In Sec. V, we relax the single-energy approximation and assess whether the energy dependence further affects the development of flavor instabilities. Finally, we discuss our findings and conclude in Sec. VI. In addition, Appendix A provides details on the correlations between neutrino creation and annihilation operators in the presence of neutrino degeneracy.
@@ II Neutrino equations of motion
$\rho_{\mathbf{p}}$ $=\begin{pmatrix}\rho_{ee;\mathbf{p}}&\rho_{ex;,\mathbf{p}}\\ \rho_{ex;\mathbf{p}}^{*}&\rho_{xx;\mathbf{p}}\end{pmatrix}\ ,$ (1a)
$\bar{\rho}_{\mathbf{p}}$ $=\begin{pmatrix}\bar{\rho}_{ee;\mathbf{p}}&\bar{\rho}_{ex;\mathbf{p}}\\ \bar{\rho}_{ex;\mathbf{p}}^{*}&\bar{\rho}_{xx;\mathbf{p}}\end{pmatrix}\ ,$ (1b)
where $\rho_{ll^{\prime};\mathbf{p}}$ and $\bar{\rho}_{ll^{\prime};\mathbf{p}}$ denote the number density ($l=l^{\prime}$) and flavor coherence ($l\neq l^{\prime}$). These quantities are defined as correlation functions of two creation and annihilation operators. Indeed, in a homogeneous medium, $\langle a^{\dagger}_{l^{\prime},\mathbf{p^{\prime}}}a_{l,\mathbf{p}}\rangle\equiv(2\pi)^{3}\delta(\mathbf{p}-\mathbf{p^{\prime}})\rho_{ll^{\prime};\mathbf{p}}$, where the angle brackets denote the average on the neutrino ensemble. If the properties of the ensemble are space dependent, then each matrix element can be defined locally by performing the Wigner transform of the correlation functions [5, 40, 33]. The results then represent an average over a certain volume (momentum spread) around the position $\mathbf{r}$ (momentum $\mathbf{p}$). However, such an approach carries the assumption that inhomogeneities only occur at large scales, e.g. due to temperature gradients in the medium. Short-range fluctuations would be associated with large momentum uncertainties, which would make it impossible to specify $\mathbf{p}$. Moreover, in an inhomogeneous system, correlations between neutrino momenta would not be negligible, so the mean-field approximation would not hold in the first place [41].
In this section, we first introduce the equations of motion of neutrinos within the mean-field approximation, and then modify them to take into account effects linked to neutrino degeneracy not included in the mean-field approach. We assume that the system is fully axisymmetric, so that the momentum distribution can be parametrized in terms of energy, $E=|\mathbf{p}|c$, and radial velocity $vc$. Moreover, we normalize the density matrices such that $(2\pi)^{-1}\int\mathrm{d}^{3}p\mathrm{Tr}(\rho_{\mathbf{p}})=1$ and $(2\pi)^{-1}\int\mathrm{d}^{3}p\mathrm{Tr}(\bar{\rho}_{\mathbf{p}})=\bar{n}_{\nu}/n_{\nu}$, where $n_{\nu}$ ($\bar{n}_{\nu}$) is the total (anti)neutrino number density. A factor $2\pi$ from the integral over the azimuthal angle is absorbed into the definition of $\rho_{\mathbf{p}}$ and $\bar{\rho}_{\mathbf{p}}$.
@@ II.1 Equations of motion within the mean-field approximation
Momentum correlations are not taken into account in the mean-field approximation. To first order in the Fermi constant, the evolution of the density matrices in Eq. (1) is determined by products of the form $\langle a^{\dagger}_{a}(\mathbf{p})a_{b}(\mathbf{q})\rangle\langle a^{\dagger}_{c}(\mathbf{p^{\prime}})a_{d}(\mathbf{q^{\prime}})\rangle$ (see Appendix A for details); this implies that the only contributions to the neutrino equations of motion are momentum swaps between (anti)neutrinos and annihilation and creation of neutrino-antineutrino pairs:
${{\nu}}_{a}(\mathbf{p})+{\nu}_{b}(\mathbf{q})$ $\to{\nu}_{a}(\mathbf{q})+{\nu}_{b}(\mathbf{p})\,,$ (2a)
$\nu_{a}(\mathbf{p})+\bar{\nu}_{a}(\mathbf{q})$ $\to\nu_{b}(\mathbf{p})+\bar{\nu}_{b}(\mathbf{q})\,,$ (2b)
These processes are represented by the following Hamiltonian:
@@
where ${D}_{n}\equiv\int\mathrm{d}^{3}p^{\prime}(2\pi E^{\prime 2})^{-1}L_{n}(v^{\prime})({\rho}_{\mathbf{p^{\prime}}}-\bar{\rho}_{\mathbf{p^{\prime}}})$. Here, $L_{n}$ are the Legendre polynomials, normalized such that $\int_{-1}^{+1}\mathrm{d}v^{\prime}L_{n}(v^{\prime})L_{m}(v^{\prime})=(n+\frac{1}{2})^{-1}$, and $v^{\prime}$ is the radial (anti)neutrino velocity.
Neutrinos can undergo different momentum-changing interactions. The interference among the different scattering amplitudes enhances the processes in Eqs. (2). Equations (4) represent the evolution of a homogeneous neutrino ensemble due to coherent scattering. If $a\neq b$, these amplitudes are indistinguishable from the flavor-changing amplitude due to vacuum neutrino mixing, so they can amplify the seed coherence provided by the latter–this is the mechanism responsible for collective flavor instabilities.
Recent literature has investigated whether correlations may alter the mean-field description of flavor conversion; for example, correlations could arise from incoherent scattering and damp collective oscillations due to quantum decoherence between the single-particle density matrices [42]. Moreover, neutrinos with the same flavor are indistinguishable, hence their quantum state has to be antisymmetric to particle exchange: hence, momentum correlations should inevitably appear in a coherent many-body state. When describing astrophysical environments, it is usually assumed that the system starts as a mixture of pure flavor eigenstates, i.e. the density matrices are diagonal and the ensemble is fully incoherent. This assumption is well motivated: (anti)neutrinos in these systems are emitted as highly localized wavepackets with a specific flavor, which may not overlap. On the other hand, two plane waves with the same momentum always interfere. The lack of momentum correlations is thus not fully justified.
@@ II.2 Beyond-mean-field effects linked to neutrino degeneracy
For momentum correlations to vanish, the neutrino ensemble must be homogeneous over a certain length scale. Locally, neutrinos can then be described as plane waves, so their momentum is exact. The consequence, as shown in Eq. (2) and detailed in Appendix A, is that coherent processes only involve two momentum modes. In contrast, if momentum correlations are taken into account, every neutrino contributing to the scattering amplitude can recoil to some extent. The inhomogeneity implied by such correlations also means that the momenta in Eq. (2) become approximate quantities. Hence, the interacting particles do not exactly swap their momenta, nor are they scattered precisely in the forward direction: rather, the initial and final momenta are indistinguishable within their uncertainty. This momentum exchange could modify the development of flavor coherence (see Fig. 2): neutrinos cannot recoil into states that are already occupied, so certain coherent processes could be Pauli blocked if the neutrino ensemble is degenerate. The suppression of coherent interactions would be described by momentum correlations between the interacting neutrinos and neighboring spectator neutrinos, representing the anti-symmetrization of the many-body neutrino quantum state.
Figure 2: Illustrative neutrino-neutrino interaction in the mean-field approximation (left) vs. beyond-mean-field case (right). In both cases, an electron neutrino (blue) scatters with a heavy-lepton-flavor neutrino (orange). In the mean-field scenario, neutrino plane waves swap their momenta: the process is indistinguishable from vacuum oscillations, and flavor coherence grows as a result. In the beyond-mean-field picture, there is a small recoil, so the final momenta are different, $\mathbf{q^{\prime}\sim p}$ and $\mathbf{p^{\prime}\sim q}$. As a result, the incoming $\nu_{e}$ wavepacket is scattered into a region where there already are $\nu_{x}$’s with similar momentum ($\mathbf{k\sim p^{\prime}}$), which do not participate in the scattering. This interaction cannot affect flavor coherence, because $\nu_{e}(\mathbf{p^{\prime}}\sim\mathbf{k})$ and $\nu_{x}(\mathbf{k})$ have to form a singlet state. We represent the neutrinos as flavor eigenstates for clarity, but in an astrophysical environment, they would be in flavor superposition due to vacuum mixing: hence, all of the particles that we sketch would be indistinguishable.
@@
Working out the form of this non-Hermitian Hamiltonian exceeds the scope of this paper. Instead, we design a toy model encapsulating beyond-mean-field effects.
The occurrence of fast instabilities depends on the presence of a crossing in the ELN angular distribution [15, 44]. However, vacuum oscillations can relax this requisite: the mass difference term in the neutrino dispersion relation can be absorbed into the definition of an effective ELN angular distribution, with stability being determined by crossings in the latter [45]. Likewise, we explore the possibility that beyond-mean-field effects in neutrino self-interactions can be described with effective density matrices, $\tilde{\rho}_{\mathbf{p}}$ and $\tilde{\bar{\rho}}_{\mathbf{p}}$.
In the many-body approach, e.g. if neutrino self-interactions were modeled through the scattering between localized wavepackets, flavor evolution should be sensitive to the momentum uncertainty of each interacting wavepacket and the spectator. Each neutrino would occupy a finite phase-space volume: thus, if an (anti)neutrino recoiled into the region occupied by an identical spectator neutrino, the overlapping wavepackets would be suppressed due to the exclusion principle. In a homogeneous single-particle description, quantum uncertainties are not defined; the matrix elements $\tilde{\rho}_{ij;\mathbf{p}}$ and $\tilde{\bar{\rho}}_{ij;\mathbf{p}}$ can only depend on $\rho_{ll^{\prime};\mathbf{q}}$ and $\bar{\rho}_{ll^{\prime};\mathbf{q}}$. The following ansatz captures our main intuition that neutrino self-interaction cannot generate flavor coherence, if the (anti)neutrino phase space is fully occupied:
@@
where $\mu_{\nu_{l}}$ is the chemical potential of neutrinos with flavor $l$, and $T$ is the local temperature.
The form of the blocking factors $(1-B_{l;\mathbf{p}})^{n}$ is motivated by the effect of Pauli blocking on incoherent processes: scattering rates are suppressed by $\prod_{i}(1-B_{i})$, where $i$ runs over all particles in the final state of the interaction. These factors come from the anticommutation relation of the neutrino creation and annihilation operators, $\{a^{\dagger}_{l,\mathbf{p}}a_{l^{\prime},\mathbf{p^{\prime}}}\}=(2\pi)^{3}\delta_{ll^{\prime}}\delta(\mathbf{p}-\mathbf{p^{\prime}})$ [5]. We choose the exponent $n=1/2$ because it is not cross sections that enter the calculation of the refractive term, but scattering amplitudes.
In the rest of the paper, we investigate the effect of such beyond-mean-field corrections due to neutrino degeneracy. We stress that while the mean-field equations presented in Ref. [5] do take into account Pauli blocking effects in neutrino-matter collisions, momentum correlations are not considered in the mean-field approach.